OpenStudy (anonymous):

Can I simplify; x=21/√3

OpenStudy (anonymous):

firstly rationalize by multiplying by \[\sqrt{3}\] both numerator and denominator

OpenStudy (anonymous):

you get \[21\sqrt{3}/3\]

OpenStudy (anonymous):

You can't simplify it. The square root of 3 is not a multiple of 21.

OpenStudy (anonymous):

to simplify rationalization is required

OpenStudy (anonymous):

you get \[7\sqrt{3}\]

OpenStudy (anonymous):

so my answer is that

OpenStudy (anonymous):

Depends, are you in algebra?

OpenStudy (anonymous):

If you are, you haven't learned rationalization. That means you can't simplify the fraction.

OpenStudy (anonymous):

can that be an answer in finding exact value of x in a triangle that have degrees of 21 & 60

OpenStudy (campbell_st):

well you seem to have a problem with you triangle... and its non right angled... so the tan ratio doesn't work.... which was the original question I answered... you you have angles of 21 and 60 the 3rd angle is 99 degrees so non right angled

OpenStudy (anonymous):